# MemControl: Mitigating Memorization in Medical Diffusion Models via Automated Parameter Selection

### v1

@@ Abstract
Diffusion models show remarkable ability in generating images that closely mirror the training distribution. However, these models are prone to training data memorization, leading to significant privacy, ethical, and legal concerns, particularly in sensitive fields such as medical imaging. We hypothesize that memorization is driven by the overparameterization of deep models, suggesting that regularizing model capacity during fine-tuning could be an effective mitigation strategy. Parameter-efficient fine-tuning (PEFT) methods offer a promising approach to capacity control by selectively updating specific parameters. However, finding the optimal subset of learnable parameters that balances generation quality and memorization remains elusive. To address this challenge, we propose a bi-level optimization framework that guides automated parameter selection by utilizing memorization and generation quality metrics as rewards. Our framework successfully identifies the optimal parameter set to be updated to satisfy the generation-memorization tradeoff. We perform our experiments for the specific task of medical image generation and outperform existing state-of-the-art training-time mitigation strategies by fine-tuning as few as 0.019% of model parameters. Furthermore, we show that the strategies learned through our framework are transferable across different datasets and domains. Our proposed framework is scalable to large datasets and agnostic to the choice of reward functions. Finally, we show that our framework can be combined with existing approaches for further memorization mitigation.
@@ 1 Introduction
Diffusion models [1, 2, 3, 4] have demonstrated exceptional proficiency in generating high-quality data across various modalities, including images [5, 6, 7, 8], audio [9, 10] and graphs [11, 12]. They underpin the commercial success of popular products such as Midjourney, Stable Diffusion and DALL-E, which make it possible to generate images based on user queries. The appeal of generative diffusion models lies in their ability to produce high-quality novel images that capture the underlying intricacies of real data. These models have diverse practical applications, for example they can be used for data augmentation that helps train more accurate recognition systems [13]. They can also be helpful for academic research, for example to help with understanding disease progression [14]. One particularly important and potentially transformative use case of synthetic data is for surpassing restrictions on sharing sensitive data.
Generative diffusion models or the generated synthetic data can be shared instead of the original data, addressesing privacy concerns while making it possible to drive progress. However, the utility of this approach would be nullified if diffusion models were to inadvertently disclose part of the underlying training data by generating images that are their near-replicas. This phenomenon is known as memorization and has linked diffusion models with various legal and ethical issues [15]. It has singled them out as potentially the least private form of generative models [16].
@@
Privacy risks related to memorization are especially significant in the context of healthcare applications. The risks include exact replication of images, potentially disclosing information about the patient [17] and making it possible to conduct re-identification attacks [18]. In re-identification attacks, models that can identify whether two images belong to the same patient can allow an attacker with partial information to link a memorized synthetic image to the associated patient. The growing concern has inspired the development of methods to mitigate memorization. These methods include training-time interventions such as de-duplication [19, 20] and noise regularizers [21], as well as inference-time mitigations such as token-rewriting [20, 22]. However, a key limitation of the fast and effective token-rewriting methods is their vulnerability to white-box attacks. An attacker with access to the model can simply disable the token-rewriting during inference to extract private data.
@@
In this paper, we introduce a new training-time intervention aimed at addressing memorization in diffusion models. We hypothesize that the overcapacity of large neural networks facilitates memorization, and thus we explore whether regularizing model capacity can mitigate this issue. In domain-specific diffusion models, such as those used for medical imaging, the typical workflow involves starting with a general-purpose diffusion model and fine-tuning it with task-specific data [23]. We posit that there is a trade-off between the extent of fine-tuning and the potential for the model to memorize in-domain medical images. If no fine-tuning is performed, the pre-trained model may not replicate specific in-domain data but will also struggle to generate high-quality domain-specific images. Conversely, extensive fine-tuning will push the model to generate high-quality in-domain images but also increase memorization risk. We investigate whether it is possible to strike an effective balance, constraining the model enough to produce high-quality generations while minimizing memorization.
@@
The key contributions of our work are summarised as follows: (1) Empirically demonstrating that reducing model capacity significantly mitigates memorization; (2) Highlighting the delicate balance between memorization and generalization and how it is influenced by model capacity (Sec. 4.2); (3) Proposing a bi-level optimization framework that automates parameter (capacity) selection to optimize both generative quality and memorization mitigation (Sec. 4.3); and (4) Showcasing that the parameter subsets discovered are transferable across different datasets and domains (Sec. 4.4).
@@ 3 Methodology
3.1 Diffusion Model Background
We now summarize the fundamentals of diffusion models. Diffusion models use a forward and reverse diffusion process. A forward diffusion process involves a fixed Markov chain spanning $T$ steps, where each step introduces a pre-determined amount of Gaussian noise. The noise is iteratively added to a data point $x_{0}$ drawn from the real data distribution $q(x_{0})$ as described by:
@@ 3.2 Memorization and Image-Quality Metrics
Our goal is to learn the optimal parameter subset for fine-tuning, given a PEFT method to limit the amount of memorization while maintaining generated image quality. There are various metrics used in the community to measure both memorization [16, 21] and generation quality [48]. In this section, we introduce the metrics that we use in this paper.
Memorization: Nearest Neighbour Distance The simplest definition of memorization considers an example $\hat{x}$ drawn from the model to be memorized if there is a training example $x$ such that $l(x,\hat{x})<\delta$ for some distance metric $l$ and threshold $\delta$. This means a training example has been memorized if the model produces a near-replica as a sample. A natural metric for the overall degree of memorization of a model is the average nearest neighbour distance between each synthetic sample and its closest example in the training set. For samples $\{\hat{x}_{i}\}$ drawn from the model and a training set $\{x_{j}\}$, we follow [49] and quantify memorization as the average minimum cosine distance (AMD):
@@
Memorization: Extraction Attack We also evaluate memorization using an image extraction attack described in [16]. The attack involves two main steps: (1) Generating numerous image samples for a given prompt, and (2) Conducting membership inference to identify memorized images. In [16], membership inference is performed using a heuristic that constructs a graph of similar samples and identifies sufficiently large cliques—sets of samples that are all similar to each other. Samples from these cliques are likely to be memorized examples. Similarity is measured using a modified $\ell_{2}$ distance, defined as the maximum $\ell_{2}$ distance across corresponding tiles of the two compared images. A smaller number of extracted images translates to lesser memorization.
@@ 3.4 Our Method: Optimizing PEFT for Mitigating Memorization
Our method employs a multi-objective bi-level optimization formulation. It trains the model in the inner loop assuming a particular PEFT mask that specifies which diffusion model parameters to freeze or learn. The PEFT mask is optimized in the outer loop so as to obtain a diffusion model with low memorisation and high generation quality. This process is formalized using Eq. 7 and Alg. 1.
@@ 4.1 Experimental Setup
Datasets: We use the MIMIC dataset [30] that consists of chest X-rays and associated radiology text reports. The dataset was split into train, validation and test splits ensuring no leakage of subjects, resulting in 110K training samples, 14K validation and 14K test samples. Note that that several images in this dataset can be associated with the same text prompt due to similarity in clinical findings.
Efficient HPO with a Memorization Subset: Conducting mask search on the whole unaltered training set is infeasible because only a minority of images are memorised, and calculating the memorization metric (Eq. 6, Sec. 3.2) is therefore noisy and slow. To alleviate this, we construct a specific training set $\mathcal{D}^{train}$ for HPO mask search consisting of (1) 1% of the original training set, and (2) a special memorization set $\mathcal{D}^{mem}$ to drive mask search. Specifically, we select 100 text-image pairs from the training set and duplicate them 50 times. With this setup the HPO training set is small, and we know which images are expected to be memorised. HPO mask search is then driven by the two validation objectives $d^{mem}$ and $d^{fid}$ (Eq. 7, Sec. 3.2) which are evaluated on the memorization set $\mathcal{D}^{mem}$ and the validation set $\mathcal{D}^{val}$ respectively. A detailed explanation on memorization subset creation is provided in Appendix C.1.
Experimental Settings: All experiments fine-tune a pre-trained Stable Diffusion model (v1.5) for 20,000 steps with a batch size of 128, using the AdamW optimizer [59]. We employ no warmup steps and utilize a constant learning rate scheduler. Hyperparameter optimization (HPO) search was implemented using Optuna package [58]. The final generation quality, indicated by the FID score ($d^{fid}$) is reported on a held-out test set, and memorization, indicated by the AMD score ($d^{amd})$, is measured on the training set. The BioViL-T score is measured between the original text captions and generated images from the test set. The extraction attack is conducted between synthetic images and the training set images.
@@ 4.2 PEFT Mask Choices Explore the Memorization-Generation Tradeoff
We first explore the influence of model capacity on memorization vs. generalization quality tradeoff by recording the FID score ($d^{fid}$) and the memorization metric ($d^{mem}$) for different parameter subsets explored during the HPO search. Each HPO iteration considers a different subset of parameters to freeze or update. The results in Figure 2 show the combined pareto front from searching within the space of SV-Diff [28], DiffFit [29] and Attention Tuning [40] PEFT options. We also report the performance of the defacto full fine-tuning [46, 14], as well as the vanilla configurations of the PEFT methods. From the performance scatter plot, we see that certain configurations favour generation quality ($d^{fid}\downarrow$), others might favour reducing memorization ($d^{mem}\downarrow$) while some can perform poorly on both fronts ($d^{fid}$uparrow, $d^{mem}\uparrow$). However, our framework also discovers the Pareto front of configurations that dominate others ($d^{fid}$downarrow, $d^{mem}\downarrow$), including configurations that are near optimal for both (bottom left). Importantly, the pareto-front excludes configurations that include the defacto full fine-tuning and PEFT variants.
Figure 2: Plot illustrating how model capacity affects the memorization vs. generation quality tradeoff. The Hyperparameter Optimization (HPO) search explored various combinations of parameter subsets during fine-tuning. Each combination results in a different model capacity, generation quality ($d^{fid}\downarrow$) and memorization ($d^{mem},\downarrow$). The performance of these subsets is compared to two default PEFT configurations (orange squares) and full fine-tuning (green square). The pareto front of optimal parameter subsets, indicated by the lowest $d^{fid}$ and $d^{mem}$, are marked in red.
@@
Given the optimal PEFT mask discovered by our framework in Sec. 4.2, we fine-tune using this mask on the full training set, and compare this against competitors. Recall that we consider two types of competitors for memorization mitigation: purpose designed regularisers [20, 22], and alternative PEFT methods [28, 29]. These mitigation methods are independent and potentially complementary, so we also go beyond prior work in reporting all their combinations as well. From the results in Table 1, we observe that: (1) Surprisingly, and despite not being designed to tackle memorization, PEFT methods SV-DIFF [28] and DiffFit [29] outperform state-of-the-art purpose designed mitigation methods such as [22] on several metrics. (2) Our MemControl framework achieves the best performance across all metrics. It outperforms both purpose designed mitigation methods, PEFT methods, and their combination across all generation quality and memorization metrics.
Table 1: Comparitive evaluation of generation quality (FID $\downarrow$, BioViL-T Score $\uparrow$) and memorization (AMD $\uparrow$, Extracted Images $\downarrow$) on the MIMIC dataset. Our MemControl outperforms purpose designed memorization mitigations, standard PEFT methods, and their combinations.
@@ 4.4 PEFT Strategy is Transferable
FT Method + Mitigation Strategy FID Score ($\downarrow$) AMD ($\uparrow$)
Freeze 62.16 0.101
Full FT 11.01 0.044
Full FT + RWA 08.72 0.048
Full FT + Threshold Mitigation 09.11 0.052
SV-DIFF 08.10 0.058
SV-DIFF + RWA 13.53 0.061
SV-DIFF + Threshold Mitigation 09.38 0.065
DiffFit 07.40 0.054
DiffFit + RWA 11.94 0.058
DiffFit + Threshold Mitigation 12.03 0.062
Attention Tuning 08.04 0.047
Attention Tuning + RWA 09.12 0.049
Attention Tuning + Threshold Mitigation 09.11 0.051
MemControl (ours) 06.34 0.093

### v2

# MemControl: Mitigating Memorization in Diffusion Models via Automated Parameter Selection

@@ Abstract
Diffusion models excel in generating images that closely resemble their training data but are also susceptible to data memorization, raising privacy, ethical, and legal concerns, particularly in sensitive domains such as medical imaging. We hypothesize that this memorization stems from the overparameterization of deep models and propose that regularizing model capacity during fine-tuning can mitigate this issue. Firstly, we empirically show that regulating the model capacity via Parameter-efficient fine-tuning (PEFT) mitigates memorization to some extent, however, it further requires the identification of the exact parameter subsets to be fine-tuned for high-quality generation. To identify these subsets, we introduce a bi-level optimization framework, MemControl, that automates parameter selection using memorization and generation quality metrics as rewards during fine-tuning. The parameter subsets discovered through MemControl achieve a superior tradeoff between generation quality and memorization. For the task of medical image generation, our approach outperforms existing state-of-the-art memorization mitigation strategies by fine-tuning as few as 0.019% of model parameters. Moreover, we demonstrate that the discovered parameter subsets are transferable to non-medical domains. Our framework is scalable to large datasets, agnostic to reward functions, and can be integrated with existing approaches for further memorization mitigation. To the best of our knowledge, this is the first study to empirically evaluate memorization in medical images and propose a targeted yet universal mitigation strategy. The code is available at this https URL.
Figure 1: Conventional full fine-tuning results in the generation of nearly identical images across different seeds for the same text prompt. Memorization is evident through the replication of artifacts (red squares), which notably occurs with high precision and is consistently observed in the generated images. This replication can lead to patient information leakage and potential re-identification. Column 1 displays the original training images, while columns 2-5 show the closest generated samples across different seeds.
@@ 1 Introduction
Diffusion models [50, 53, 23, 54] have proven highly effective in generating high-quality data across various modalities, including images [16, 32, 43, 45], audio [33, 29], and graphs [64, 58]. They power popular products like Midjourney, Stable Diffusion, and DALL-E, enabling image generation based on user prompts. The appeal of these models lies in their ability to create high-quality novel images that capture real-world complexities, with applications ranging from data augmentation for more accurate recognition systems [57] to academic research on disease progression [36]. A key use case is sharing synthetic data to circumvent privacy restrictions, driving progress while addressing privacy concerns. However, this utility is compromised if diffusion models inadvertently disclose training data by generating near-replicas—a phenomenon known as memorization, which raises significant legal and ethical concerns [7] and identifies diffusion models as potentially the least private form of generative models [8].
Significance in Medical Image Analysis:  Privacy risks related to memorization are especially significant in the context of healthcare applications. The risks include exact replication of images, potentially disclosing information about the patient [65] and making it possible to conduct re-identification attacks [20]. In re-identification attacks, models that can identify whether two images belong to the same patient can allow an attacker with partial information to link a memorized synthetic image to the associated patient. Despite of the significant threat that memorization poses in the medical domain, very limited number of studies have advocated for this issue [20, 30, 1, 14]. Furthermore, to the best of our knowledge, none of the previous works have investigated the efficacy of existing training time mitigation strategies. To bridge this gap, we focus our work on the task of medical image generation and introduce a new training-time memorization intervention method.
@@
The growing concern for training data memorization has inspired the development of methods to mitigate memorization. These methods include training-time interventions such as de-duplication [61, 52] and noise regularizers [51], as well as inference-time mitigations such as token-rewriting [52, 62]. Firstly, de-duplication is not possible in the specific case of medical images due to high-similarity between samples. Furthermore, a key limitation of the fast and effective token-rewriting methods is their vulnerability to white-box attacks. An attacker with access to the model can simply disable the token-rewriting during inference to extract private data.
In this paper, we introduce a novel training-time intervention to address memorization in diffusion models. We hypothesize that the overcapacity of large neural networks contributes to memorization and explore whether regularizing model capacity can mitigate this issue. While similar hypotheses have been investigated in previous work, our approach differs in terms of the task (text-to-image generation as opposed to language generation [24, 55], image classification [40]), the mechanism for regulating capacity (PEFT as opposed to model compression [24], localization [40, 55, 11]), and the fine-tuning paradigm employed [21].
@@
In domain-specific diffusion models, such as those used for medical imaging, the typical workflow involves starting with a general-purpose diffusion model and fine-tuning it with task-specific data [10]. We posit that there is a trade-off between the extent of fine-tuning and the potential for the model to memorize in-domain medical images. If no fine-tuning is performed, the pre-trained model may not replicate specific in-domain data but will also struggle to generate high-quality domain-specific images. Conversely, extensive fine-tuning will push the model to generate high-quality in-domain images but also increase memorization risk. We investigate whether it is possible to strike an effective balance, constraining the model enough to produce high-quality generations while minimizing memorization.
@@
The key contributions of our work are summarised as follows: (1) We conduct the first study that empirically evaluates memorization mitigation mechanisms for medical images, establishing a fair benchmark for future studies; (2) Empirically demonstrating that reducing model capacity significantly mitigates memorization; (3) Highlighting the delicate balance between memorization and generalization and how it is influenced by model capacity (Sec. 4.2); (4) Proposing a bi-level optimization framework that automates parameter (capacity) selection to optimize both generative quality and memorization mitigation (Sec. 4.3); and (5) Showcasing that the parameter subsets discovered are transferable across different datasets and domains (Sec. 4.4).
@@ 3 Methodology
3.1 Preliminaries
Diffusion models use a forward and reverse diffusion process. A forward diffusion process involves a fixed Markov chain spanning $T$ steps, where each step introduces a pre-determined amount of Gaussian noise. The noise is iteratively added to a data point $x_{0}$ drawn from the real data distribution $q(x_{0})$ as described by:
@@ 3.2 Memorization and Image-Quality Metrics
Our goal is to learn the optimal parameter subset for fine-tuning, given a PEFT method to limit the amount of memorization while maintaining generated image quality. There are various metrics used in the community to measure both memorization [8, 51] and generation quality [4], and we introduce the metrics that we use in this paper next.
Memorization: Nearest Neighbour Distance The simplest definition of memorization considers an example $\hat{x}$ drawn from the model to be memorized if there is a training example $x$ such that $l(x,\hat{x})<\delta$ for some distance metric $l$ and threshold $\delta$. This means a training example has been memorized if the model produces a near-replica as a sample. A natural metric for the overall degree of memorization of a model is the average nearest neighbour distance between each synthetic sample and its closest example in the training set. For $N$ samples $\{\hat{x}_{i}\}$ drawn from the model and a training set $\{x_{j}\}$, we follow [3] and quantify memorization as the average minimum cosine distance (AMD):
@@
Memorization: Extraction Attack We also evaluate memorization using an image extraction attack described in [8]. The attack involves two main steps: (1) Generating numerous image samples for a given prompt, and (2) Conducting membership inference to identify memorized images. In [8], membership inference is performed using a heuristic that constructs a graph of similar samples and identifies sufficiently large cliques—sets of samples that are all similar to each other. Samples from these cliques are likely to be memorized examples. Similarity is measured using a modified $\ell_{2}$ distance, defined as the maximum $\ell_{2}$ distance across corresponding tiles of the two compared images. A smaller number of extracted images translates to less memorization, and we measure it during evaluation.
@@ 3.4 Our Method: Optimizing PEFT for Mitigating Memorization
Our method employs a multi-objective bi-level optimization formulation. It trains the model in the inner loop assuming a particular PEFT mask that specifies which diffusion model parameters to freeze or learn. The PEFT mask is optimized in the outer loop so as to obtain a diffusion model with low memorisation and high generation quality. This process is formalized using Eq. 7 and Algorithm 1:
@@
such $\text{such that}\quad\Delta\phi^{*}=\operatorname*{arg\,min}_{\Delta\phi}\mathcal{L}^{base}\left(\mathcal{D}^{train};\theta_{0}+\omega\odot\Delta\phi\right),$
@@
Additional details on HPO sampling and a detailed analysis of the scalability and convergence of the HPO are presented in the supplementary material (Section 4).
@@ 4.1 Experimental Setup
Datasets: We use the MIMIC dataset [31] that consists of chest X-rays and associated radiology text reports. The MIMIC dataset, because of its scale and rich annotations stands as the standard dataset for performing text-to-image generation in medical image analysis [10, 18]. The dataset was split into train, validation and test splits ensuring no leakage of subjects, resulting in 110K training samples, 14K validation and 14K test samples. Note that that several images in this dataset can be associated with the same text prompt due to similarity in clinical findings.
Efficient HPO with a Memorization Subset: Searching for the optimal parameter subset (mask) across the entire training set is computationally infeasible. Additionally, only a small fraction of images within the training set are actually memorized. Consequently, calculating the memorization metric (Eq. 6, Sec. 3.2) would be both noisy and slow if conducted on the entire dataset. To address this, we construct a specific training set, $\mathcal{D}^{train}$, for mask search during the HPO, comprising of (1) 1% of the original training set and (2) a specialized memorization set, $\mathcal{D}^{mem}$.
Creation of $\mathcal{D}^{mem}$: We select 100 image-text pairs from the training set and duplicate each 50 times, following the approach in [62, 44]. This duplication forces the model to memorize these specific samples during fine-tuning, providing a clear indication of which pairs are memorized. Unlike natural imaging datasets such as LAION [49], where memorized prompts have been identified in prior studies [61], no such information exists for medical datasets like MIMIC. Therefore, we explicitly create a subset for this task. We follow a specific heuristic for selecting the 100 the image-text pairs for $\mathcal{D}^{mem}$, detailed in Supplementary material (Section 2). Following this setup, the HPO training set is small, and we know which images are expected to be memorized. HPO mask search is then driven by the two validation objectives $d^{mem}$ and $d^{fid}$ (Eq. 7, Sec. 3.2) which are evaluated on the memorization set $\mathcal{D}^{mem}$ and the validation set $\mathcal{D}^{val}$ respectively.
Experimental Settings: All experiments fine-tune a pre-trained Stable Diffusion model (v1.5) for 20,000 steps with a batch size of 128, using the AdamW optimizer [39]. The HPO search was implemented using Optuna package [2]. The FID score ($d^{fid}$) is reported on a held-out test set, and memorization, indicated by the AMD score ($d^{amd})$, is measured on the training set. The BioViL-T score is measured between the original text captions and generated images from the test set. The extraction attack is conducted between synthetic images and the training set images.
@@ 4.2 PEFT Mask Choices Explore the Memorization-Generation Tradeoff
We first explore the influence of model capacity on memorization vs. generalization quality tradeoff by recording the FID score ($d^{fid}$) and the memorization metric ($d^{mem}$) for different parameter subsets explored during the HPO search. Each HPO iteration considers a different subset of parameters to freeze or update. The results in Figure 3 show the combined pareto front from searching within the space of SV-Diff [22], DiffFit [63] and Attention Tuning [56] PEFT options. We also report the performance of the defacto full fine-tuning [18, 36], as well as the vanilla configurations of the PEFT methods.
From the performance scatter plot, we see that certain configurations favour generation quality ($d^{fid}\downarrow$), others might favour reducing memorization ($d^{mem}\downarrow$) while some can perform poorly on both fronts ($d^{fid}$uparrow, $d^{mem}\uparrow$). However, our framework also discovers the Pareto front of configurations that dominate others ($d^{fid}$downarrow, $d^{mem}\downarrow$), including configurations that are near optimal for both (bottom left). Importantly, the pareto-front excludes configurations that include the defacto full fine-tuning and PEFT variants.
Figure 3: Plot illustrating how model capacity affects the memorization vs. generation quality tradeoff. The HPO search explored various combinations of parameter subsets during fine-tuning (blue markers). Each combination results in a different model capacity, generation quality ($d^{fid}\downarrow$) and memorization ($d^{mem},\downarrow$). The performance of these subsets is compared to two default PEFT configurations (orange squares) and full fine-tuning (green square). The pareto front of optimal parameter subsets, indicated by the lowest $d^{fid}$ and $d^{mem}$, are marked in red, while the final optimal solution is marked in cyan.
@@
Given the optimal PEFT mask discovered by our framework in Sec. 4.2, we fine-tune using this mask on the full training set, and compare this against competitors. Recall that we consider two types of competitors for memorization mitigation: purpose designed regularisers [52, 62], and alternative PEFT methods [22, 63]. These mitigation methods are independent and potentially complementary, so we also go beyond prior work in reporting all their combinations as well. Quantitatively, the results in Table 1 show that: (1) Despite not being designed to tackle memorization, PEFT methods SV-DIFF [22] and DiffFit [63] outperform state-of-the-art purpose designed mitigation methods such as [62] on several metrics. (2) Our MemControl framework achieves the best performance across all metrics. It outperforms both purpose designed mitigation methods, PEFT methods, and their combination across all generation quality and memorization metrics. As a qualitative comparison, we show the synthetic generations in Figure 4. Firstly, it can be directly observed that samples generated with full fine-tuning are near-identical replicas of the ground-truth training images, also highlighted in Figure 1. Additionally, combining full fine-tuning with a mitigation strategy (RWA) negatively impacts the generation performance. Finally, MemControl enables high-quality generation and prevents replication simultaneously.
@@
MemControl + RWA [52] 013.10 0.121 00.69 023
MemControl + Threshold Mitigation [62] 011.31 0.142 00.75 016
Table 1: Comparative evaluation of generation quality (FID $\downarrow$, BioViL-T Score $\uparrow$) and memorization (AMD $\uparrow$, Extracted Images $\downarrow$) on the MIMIC dataset. Our MemControl outperforms purpose-designed memorization mitigations, standard PEFT methods, and their combinations.
Additionally, MemControl can be seamlessly integrated with these approaches and effectively complement existing methods to enhance memorization mitigation (bottom two rows).
@@ 4.4 PEFT Strategy is Transferable
FT Method + Mitigation Strategy FID Score ($\downarrow$) AMD ($\uparrow$) Num. Extracted Images ($\downarrow$)
Freeze 62.16 0.101 105
Full FT 11.01 0.044 226
Full FT + RWA 08.72 0.048 198
Full FT + Threshold Mitigation 09.11 0.052 181
SV-DIFF 08.10 0.058 172
SV-DIFF + RWA 13.53 0.061 163
SV-DIFF + Threshold Mitigation 09.38 0.065 158
DiffFit 07.40 0.054 144
DiffFit + RWA 11.94 0.058 177
DiffFit + Threshold Mitigation 12.03 0.062 135
Attention Tuning 08.04 0.047 194
Attention Tuning + RWA 09.12 0.049 192
Attention Tuning + Threshold Mitigation 09.11 0.051 182
MemControl (ours) 06.34 0.093 108